# On-chip multi-timescale spatiotemporal optical synchronization

### v2

@@ VI Summary and outlook
Here for the first time, we demonstrated topological spatiotemporal mode-locking in a multi-timescale 2D array of hundreds of coupled ring resonators. Our results indicate topological mode-locking and non-harmonic multi-timescale synchronization between photonic modes of hundreds of optical resonators is possible in a compact chip architecture. Moreover, we demonstrated clear spatial, optical, and electrical noise signatures for the identification and classification of different types of edge and bulk states, which constitutes a way forward for understanding complex nonlinear states of light in topological photonics. Going forward, the uncoupled nature of the two timescales in our mode-locking scheme opens the possibility of independently tuning the timescales into a variety of configurations. Moreover, the nesting in our system can be cascaded further by adding another lattice layer with an independent coupling rate, opening the door for non-harmonic many-timescale mode-locking from MHz to THz, as well as exploring nonlinear multi-mode optical synchronization phenomena [20].
@@
On a more fundamental level, our results provide a new platform to study the interplay of topology and optical nonlinearities [32, 33, 34, 25], as well as intriguing nonlinear topological physics unique to bosons [35, 32], such as nonlinearity-induced restructuring of the bulk-edge correspondence [36, 37] and nonlinear control of topology [38, 36, 39, 40]. Looking forward, beyond our AQH lattice, intriguing avenues open up to explore nonlinear optical synchronization phenomena in the presence of other synthetic gauge fields, such as Floquet [41] and integer Hall models, other topological models [42, 43] and frequency regimes [44], and Kerr-induced symmetry-breaking effects that are unique to coupled-resonator systems [45, 46, 23]. Moreover, the nested character of our combs provides a new and higher-dimensional testbed for exploring comb spectroscopy over different precision scales, as well as fundamentals of noise physics in coherent solitonic solutions [27]. Further, our platform may be adapted to study topology in other mode-locked photonic sources, including synchronously pumped OPOs [47, 48, 49], topological temporally mode-locked lasers [50], and dissipative Kerr cavities [51]. The gradual tunability of our edge-to-bulk combs also can facilitate the investigation of theoretically proposed chimera-like states -coexisting chaotic and coherent states of the electromagnetic field inside the cavity [52].
VII Methods
VII.1 Device fabrication
@@
The topological photonic devices were fabricated in a commercial foundry in the same way as the ones presented in Ref. [26]. A high-resolution optical image presented in Figure S1 shows the topological photonic lattice used in this work. The microring resonators are made of Si3N4 embedded in SiO2. Their dimensions are 1200 nm wide and 800 nm thick. The gaps between the microrings, as well as those between the bus waveguides and the microrings, are 300 nm.
VII.2 Experimental setup
@@
For the linear measurements as well as the nonlinear imaging, we closely followed the method presented in Ref. [26].
@@
VIII Acknowledgements
The authors wish to acknowledge fruitful discussions with Apurva Padhye, Deric Session, Avik Dutt, and Daniel Leykam. This work was supported by ONR N00014-20-1-2325, and MURI FA9550-22-1-0339.
IX Authors contribution
@@ X Competing interests
XI Data availability
All of the data that support the findings of this study are reported in the main text and Supplementary Information. Source data are available from the corresponding authors on reasonable request.
XII Supplementary Materials for
@@
XIII S1: Design and Hamiltonian description
@@
The device design and Hamiltonian description of our system, including the linear and nonlinear dynamics as well as drive and dissipation closely follow the device described in Ref. [26].
XIV S2: Device parameter estimation
@@
A high-resolution optical image in 5c illustrates the topological photonic lattice utilized in this study. The device comprises a two-dimensional array of 261 interconnected photonic ring resonators with a pair of coupled input-output waveguides. These waveguides, embedded in silicon dioxide, have dimensions of 1200 nm in width and 800 nm in thickness to ensure operation within the anomalous dispersion regime. Simulated mode profiles and dispersion characteristics are presented in Figure S2 of the Supplementary Material of Ref. [53]. Each ring features a racetrack structure, incorporating m straight coupling sections and Euler bend regions, with an effective radius of m, resulting in a free spectral range (FSR) of approximately THz.
@@
The coupling gaps between the resonators, as well as those between the input-output waveguides and the resonators, are set to 300 nm, yielding an approximate coupling strength, , of GHz. The extrinsic and intrinsic couplings, denoted as and , are estimated at GHz and GHz, respectively (see Figure S3). Further details on these calculations can be found in the Supplementary Material of Ref. [26].
XV S3: Band structure
@@
Figure 5a-b presents the unit cell and the calculated band structure of a semi-infinite AQH lattice, periodic along one axis and finite along the other. A -wide edge band region, marked in green, lies between two bulk bands marked in gray. In contrast to the bulk modes, which do not possess well-defined momentum, two unidirectional edge bands with opposite pseudospins exist. These edge states propagate in opposite directions and exhibit robustness against local disorder [29, 30].
XVI S4: Linear characterization
@@
This section closely follows the method described in Ref. [26]. A high-resolution optical image of the AQH lattice is shown in Figure 5c. The linear characterization (using weak continuous-wave tunable lasers) of the device is shown in Figure 5d-e. This includes the device’s output measured from the drop and through ports over many longitudinal (top row) as well as individual (middle row) modes, in addition to the pump wavelength range used during all the nonlinear measurements (bottom row). Moreover, Figure 5f shows the corresponding linear-regime group delay results of Figure 5d-e measured from the drop port of the device analyzed with a vector analyzer.
XVII S5: Experimental setup
@@
Figure 6 also provides a schematic of the experimental setup used for nonlinear measurements. In this setup, a pulsed tunable laser is directed into a free-space optical arrangement, which includes a variable attenuator and a polarization controller made up of a quarter-wave plate, a half-wave plate, and another quarter-wave plate. The output is subsequently coupled into a short tapered fiber and edge-coupled into the SiN chip via the input port. Coupling losses for each coupler are estimated at 2 dB to 3 dB. The combined pump and comb output is collected from the drop port using another tapered fiber and can be optionally attenuated or filtered before being analyzed. The output is then directed to broadband low-resolution (40 pm) and high-resolution (0.4 pm) optical spectrum analyzers (OSAs) based on gratings and heterodyne detection, respectively, as well as to an electrical spectrum analyzer (ESA) and an oscilloscope.
@@
Based on the linear Hamiltonian description discussed in Ref. [26], the single-mode linear simulations of the drop port transmission spectrum and corresponding experimental data of our device are shown in Figure 7a. Figure 7b-c shows the corresponding eigenvalues as well as selected spatial intensity profiles of a few eigenmodes. It can be clearly seen that the degree of confinement of the edge modes decreases by moving the pump frequency away from the center of the edge band, eventually entering the bulk band with distinct bulk profiles.
@@
Moreover, as discussed earlier, we note that in our device is designed to be 25 GHz, which leads to the broadening and merging of edge resonance within the edge band. Figure 7d shows how by reducing to smaller values (here calculated for a fixed and ), the edge modes can be spectrally resolved.
XIX S7: Nonlinear imaging
@@
Figure 8 shows the measurement details for the spatial intensity profiles of the generated frequency combs. Figure 8a shows the pump power dependence of an edge comb (pump wavelength 1547.83 nm). Since the pump is filtered and only part of the generated comb is sent to the IR camera, below the OPO threshold no light is seen. Above the OPO threshold, however, proportional to the pump power, scattered out-of-plane light can be observed.
@@
Figure 8b shows several intensity profiles for different bulk and edge combs. The strong edge confinement of the edge combs in contrast to the bulk combs can be clearly seen. The same spectral filtering that was applied to all the imaging cases is also shown schematically in Figure 8c. Moreover, highly magnified imaging of the scattered light from a single plaquette of the AQH is also shown in Figure 8d.
XX S8: Nestedness of the combs
@@
A detailed measurement of the nestedness of the topological combs is presented in Figure 9. Qualitatively, comparable nesting characteristics are observed across all the single ring modes, , which are within the bandwidth of our high-resolution OSA (1520 nm to 1630 nm). For edge combs, the comb teeth are nested. As the pump is tuned into the adjacent two bulk bands, the nesting diminishes. Moreover, the data shown for two different pump powers shows no significant change in the nestedness of the combs.
XXI S9: Optical linewidth Analysis
@@
Using the high-resolution (0.4 pm) heterodyne-based OSA, we performed detailed optical linewidth analysis of the pump laser, as well as the nested comb teeth. Figure 10a shows the pump spectrum, which has several peaks, with most of the power concentrated at its dominant primary peak. The Lorentizian fit of the primary peak reveals a FWHM of around 2.77 pm, with an estimated error of pm. Figure 10b-c shows a typical fitting procedure of the nested comb teeth in our work.
XXII S10: Pump spectral and temporal calibration
@@
We characterized the pump with the heterodyne-based OSA at several pump wavelengths as shown in Figure 11a-b. The pump was also measured directly by sending it to the PD and monitoring the RF output with the oscilloscope, which is shown in Figure 11c-d. The 250 kHz (4000 ns) repetition rate of the pump can also be seen.
XXIII S11: Fourier transform
@@
The finite duration of our pulsed pump puts a lower bound on the frequency resolution in our ESA measurements. This is shown for different scenarios in Figure 12. With our 5 ns pulse, the round-trip time of around 300 ps to 400 ps, a frequency resolution limit of around 223 MHz is calculated. We note that this limit only depends on the pulse duration, diminishing for longer pulse duration and vice versa. We also note that this limit doesn’t depend on the pulse separation as shown in Figure 12b.
XXIV S12: Notch filter
@@
In our nonlinear measurement, we filter out the pump optionally with a tunable notch filter and mostly only analyze the remaining part of the comb spectrum. The filter is chosen such that it has a consistent and very high extinction ratio over the full pump wavelength range used in our work, which is 1547.6 nm to 1548.1 nm. The full calibration of the notch filter of the pump with and without going through the device in the linear regime, as well as in the above-OPO regime is shown in Figure 13.
XXV S13: OSA, ESA and PD measurements of the combs
@@
The detailed pump power and pump wavelength dependence of the topological combs are summarized in Figure 14. The top row shows the measured comb powers as a function of pump wavelength over the 1547.6 nm to 1458.1 nm range of the nonlinear measurements. The edge combs have higher power than the bulk counterparts except for the very high pumping regime. The middle and bottom panels show the corresponding OSA and ESA spectrums. The pump was notched out in all of these measurements.
XXVI S14: Quadratic fitting of the comb optical linewidth
@@
We analyzed the optical linewidth of nested comb lines for a selected single ring mode . The results for the nested teeth 1 to 7 are shown in Figure 15 for several pump wavelengths. We clearly observe a quadratic form (with some variation) of the optical linewidth, which goes close to the pump’s linewidth near the pumped mode for most cases. Moreover, the tooth with the minimum linewidth follows the pump wavelength sweeping direction. This behavior is only observed for the edge combs, and the bulk counterparts lack such features.
@@
Moreover, Figure 16 shows the quadratic variation of the comb tooth optical linewidth over both fast and slow timescales for an edge comb. This serves as direct evidence for simultaneous mode-locking over both timescales.
@@
Further, we compare the linewidth of our topological combs with a comparable (with an identical design) single-ring race-track counterpart. Figure 17a shows a generated comb using the same pump laser in our experimental setup. As shown in Figure 17b, the optical linewidth of the comb teeth lacks the quadratic trend of the topological counterpart and also exhibits much broader linewidth compared to the pump’s linewidth. Moreover, as shown in Figure 17a, nested structure is absent in the comb teeth of the single ring comb.
XXVII S15: Temporal analysis of the pump power dependence of the bulk combs
@@
For a typical bulk comb, the temporal dynamic of the comb is measured as a function of pump power. The results are shown in Figure 18. It can be seen that no appreciable temporal patterns emerge above the OPO threshold leading to no appearance of a well-defined frequency component in the (Fourier transform) FT analysis.
XXVIII S16: Temporal analysis of the pump power dependence of the edge combs
@@
For a typical edge comb, the temporal dynamic of the comb is measured as a function of pump power. The results are shown in Figure 19. Unlike the bulk combs, it can be seen that temporal patterns emerge above the OPO threshold leading to the appearance of well-defined frequency components in the FT analysis.
@@
XXIX S17: Temporal analysis of the pump power and pump wavelength dependence of the edge and bulk combs
@@
For several edge and bulk combs, the temporal dynamic of the comb is measured as a function of pump power. The results are shown in Figure 20. The sharp contrast between edge combs and their bulk counterparts can be seen.
XXX S18: Temporal analysis of the combs as a function of pump wavelength.
@@
For a fixed pump power of 185 mW, the temporal dynamic of the combs is measured as a function of pump wavelength. The results are shown in Figure 21. The gradual emergence of temporal patterns as the pump is tuned towards the edge band can be seen.

### v3

@@ Discussion
Here for the first time, we demonstrated signatures of topological spatiotemporal mode-locking in a multi-timescale 2D array of hundreds of coupled ring resonators. We further provide results for a second copy of the device with different parts of the comb, establishing the repeatability of the experiment and strengthening our observation, as shown in SI section S20. We note that due to limitations of our experiment as shown in SI section S22, further studies are needed to confirm the existence of topological solitons (?). Nevertheless, our results indicate topological mode-locking and non-harmonic multi-timescale synchronization between photonic modes of hundreds of optical resonators is possible in a compact chip architecture. Moreover, we demonstrated clear spatial, optical, and electrical noise signatures for the identification and classification of different types of edge and bulk states, which not only adds numerous understanding to the preceding work, but also constitutes a way forward for understanding complex nonlinear states of light in topological photonics. Going forward, the uncoupled nature of the two timescales in our mode-locking scheme opens the possibility of independently tuning the timescales into a variety of configurations (?, ?). Moreover, the nesting in our system can be cascaded further by adding another lattice layer with an independent coupling rate, opening the door for non-harmonic many-timescale mode-locking from MHz to THz, as well as exploring nonlinear multi-mode optical synchronization phenomena (?).
@@
On a more fundamental level, our results provide an innovative platform to study the interplay of topology and optical nonlinearities (?, ?, ?, ?), as well as intriguing nonlinear topological physics unique to bosons (?, ?), such as nonlinearity-induced restructuring of the bulk-edge correspondence (?, ?) and nonlinear control of topology (?, ?, ?, ?). Looking forward, beyond our AQH lattice, intriguing avenues open up to explore nonlinear optical synchronization phenomena in the presence of other synthetic gauge fields, such as Floquet (?) and integer Hall models, other topological models (?, ?) and frequency regimes (?), and Kerr-induced symmetry-breaking effects that are unique to coupled-resonator systems (?, ?, ?). Moreover, the nested character of our combs provides a higher-dimensional testbed for exploring comb spectroscopy over different precision scales, as well as fundamentals of noise physics in coherent solitonic solutions (?). Further, our platform may be adapted to study topology in other mode-locked photonic sources, including synchronously pumped OPOs (?, ?, ?), topological temporally mode-locked lasers (?), and dissipative Kerr cavities (?). The gradual tunability of our edge-to-bulk combs also can facilitate the investigation of theoretically proposed chimera-like states -coexisting chaotic and coherent states of the electromagnetic field inside the cavity (?).
Materials and Methods
Device fabrication
@@
The topological photonic devices were fabricated in a commercial foundry in the same way as the ones presented in Ref. (?). A high-resolution optical image presented in Figure S1 shows the topological photonic lattice used in this work. The microring resonators are made of Si3N4 embedded in SiO2. Their dimensions are 1200 nm wide and 800 nm thick. The gaps between the microrings, as well as those between the bus waveguides and the microrings, are 300 nm.
Experimental setup
@@
For the linear measurements as well as the nonlinear imaging, we closely followed the method presented in Ref. (?).
@@
Acknowledgements
The authors wish to acknowledge fruitful discussions with Apurva Padhye, Deric Session, Avik Dutt, and Daniel Leykam.
Funding
This work was supported by ONR N00014-20-1-2325, and MURI FA9550-22-1-0339.
Authors contribution
@@ Acknowledgements
Data availability
All of the data that support the findings of this study are reported in the main text and Supplementary Information. Source data are available at https://zenodo.org/records/15492780.
Supplementary materials
Supplementary Text
Figs. S1 to S24
Movie S1
Supplementary Materials for
@@
Lida Xu1†, Mahmoud Jalali Mehrabad1∗†, Christopher J. Flower1†
Gregory Moille2, Alessandro Restelli2, Daniel G. Suarez-Forero1, Yanne Chembo3
Sunil Mittal4, Kartik Srinivasan2, Mohammad Hafezi1∗
∗Corresponding author. Email: mjalalim@umd.edu, hafezi@umd.edu
†These authors contributed equally to this work.
This PDF file includes:
Supplementary Text
Figures S1 to S24
Other Supplementary Materials for this manuscript:
Movie S1
S1: Design and Hamiltonian description
@@
The device design and Hamiltonian description of our system, including the linear and nonlinear dynamics as well as drive and dissipation, closely follow the device described in Ref. (?).
S2: Device parameter estimation
@@
A high-resolution optical image in S1c illustrates the topological photonic lattice utilized in this study. The device comprises a two-dimensional array of 261 interconnected photonic ring resonators with a pair of coupled input-output waveguides. These waveguides, embedded in silicon dioxide, have dimensions of 1200 nm in width and 800 nm in thickness to ensure operation within the anomalous dispersion regime. Simulated mode profiles and dispersion characteristics are presented in Figure S3 of the Supplementary Material of Ref. (?). Each ring features a racetrack structure, incorporating m straight coupling sections and Euler bend regions, with an effective radius of m, resulting in a free spectral range (FSR) of approximately THz.
@@
The coupling gaps between the resonators, as well as those between the input-output waveguides and the resonators, are set to 300 nm, yielding an approximate coupling strength, , of GHz. The extrinsic and intrinsic couplings, denoted as and , are estimated at GHz and GHz, respectively (see Figure S3). Further details on these calculations can be found in the Supplementary Material of Ref. (?).
The fast timescale (1 THz) comes from the length L, the material, and geometry of the single rings that construct the lattice. Consider an ideal circular ring resonator with group refractive index , denoting speed of light as , the Free Spectral Range (FSR) for a single ring is . The slow timescale corresponds to the spacing between two adjacent super-ring modes in the edge band, as shown in Figure S3. The coupling strength J between two adjacent rings is controlled by their gap. The size of the lattice is described by the number of rings N on one side. The whole edge band spans 2J, and there are N edge modes with linear dispersion inside the edge band, giving us a slow timescale . Since lattice parameters are independent of single ring parameters, the fast and slow timescales are independently designable.
S3: Band structure
@@
Figure S1a-b presents the unit cell and the calculated band structure of a semi-infinite AQH lattice, periodic along one axis and finite along the other. A -wide edge band region, marked in green, lies between two bulk bands marked in gray. In contrast to the bulk modes, which do not possess well-defined momentum, two unidirectional edge bands with opposite pseudospins exist. These edge states propagate in opposite directions and exhibit robustness against local disorder (?, ?).
S4: Linear characterization
@@
This section closely follows the method described in Ref. (?). A high-resolution optical image of the AQH lattice is shown in Figure S1c. The linear characterization (using weak continuous-wave tunable lasers) of the device is shown in Figure S1d-e. This includes the device’s output measured from the drop and through ports over many longitudinal (top row) as well as individual (middle row) modes, in addition to the pump wavelength range used during all the nonlinear measurements (bottom row). Moreover, Figure S1f shows the corresponding linear-regime group delay results of Figure S1d-e measured from the drop port of the device analyzed with a vector analyzer.
S5: Experimental setup
@@
Figure S2 also provides a schematic of the experimental setup used for nonlinear measurements. In this setup, a pulsed tunable laser is directed into a free-space optical arrangement, which includes a variable attenuator and a polarization controller made up of a quarter-wave plate, a half-wave plate, and another quarter-wave plate. The output is subsequently coupled into a short tapered fiber and edge-coupled into the SiN chip via the input port. Coupling losses for each coupler are estimated at 2 dB to 3 dB. The combined pump and comb output is collected from the drop port using another tapered fiber and can be optionally attenuated or filtered before being analyzed. The output is then directed to broadband low-resolution (40 pm) and high-resolution (0.4 pm) optical spectrum analyzers (OSAs) based on gratings and heterodyne detection, respectively, as well as to an electrical spectrum analyzer (ESA) and an oscilloscope.
@@
Based on the linear Hamiltonian description discussed in Ref. (?), the single-mode linear simulations of the drop port transmission spectrum and corresponding experimental data of our device are shown in Figure S3a. Figure S3b-c shows the corresponding eigenvalues as well as selected spatial intensity profiles of a few eigenmodes. Note that here the output ring is located upper left. It can be clearly seen that the degree of confinement of the edge modes decreases by moving the pump frequency away from the center of the edge band, eventually entering the bulk band with distinct bulk profiles.
@@
Moreover, as discussed earlier, we note that in our device is designed to be 25 GHz, which leads to the broadening and merging of edge resonance within the edge band. Figure S3d shows how by reducing to smaller values (here calculated for a fixed and ), the edge modes can be spectrally resolved.
S7: Nonlinear imaging
@@
Figure S4 shows the measurement details for the spatial intensity profiles of the generated frequency combs. Figure S4a shows the pump power dependence of an edge comb (pump wavelength 1547.83 nm). Since the pump is filtered and only part of the generated comb is sent to the IR camera, below the OPO threshold no light is seen. Above the OPO threshold, however, proportional to the pump power, scattered out-of-plane light can be observed.
@@
Figure S4b shows several intensity profiles for different bulk and edge combs. The strong edge confinement of the edge combs in contrast to the bulk combs can be clearly seen. The same spectral filtering that was applied to all the imaging cases is also shown schematically in Figure S4c. Moreover, highly magnified imaging of the scattered light from a single plaquette of the AQH is also shown in Figure S4d.
S8: Nestedness of the combs
@@
A detailed measurement of the nestedness of the topological combs is presented in Figure S5. Qualitatively, comparable nesting characteristics are observed across all the single ring modes, , which are within the bandwidth of our high-resolution OSA (1520 nm to 1630 nm). For edge combs, the comb teeth are nested. As the pump is tuned into the adjacent two bulk bands, the nesting diminishes. Moreover, the data shown for two different pump powers shows no notable change in the nestedness of the combs.
S9: Optical linewidth Analysis
@@
Using the high-resolution (0.4 pm) heterodyne-based OSA, we performed detailed optical linewidth analysis of the pump laser, as well as the nested comb teeth. Figure S6a shows the pump spectrum, which has several peaks, with most of the power concentrated at its dominant primary peak. The Lorentzian fit of the primary peak reveals a FWHM of around 2.77 pm, with an estimated error of pm. Figure S6b-c shows a typical fitting procedure of the nested comb teeth in our work.
S10: Pump spectral and temporal calibration
@@
We characterized the pump with the heterodyne-based OSA at several pump wavelengths as shown in Figure S7a-b. The pump was also measured directly by sending it to the PD and monitoring the RF output with the oscilloscope, which is shown in Figure S7c-d. The 250 kHz (4000 ns) repetition rate of the pump can also be seen.
S11: Fourier transform
@@
The finite duration of our pulsed pump puts a lower bound on the frequency resolution in our ESA measurements. This is shown for different scenarios in Figure S8. With our 5 ns pulse, the round-trip time of around 300 ps to 400 ps, a frequency resolution limit of around 223 MHz is calculated. We note that this limit only depends on the pulse duration, diminishing for longer pulse durations and vice versa. We also note that this limit doesn’t depend on the pulse separation as shown in Figure S8b.
S12: Notch filter
@@
In our nonlinear measurement, we filter out the pump optionally with a tunable notch filter and mostly only analyze the remaining part of the comb spectrum. The filter is chosen such that it has a consistent and very high extinction ratio over the full pump wavelength range used in our work, which is 1547.6 nm to 1548.1 nm. The full calibration of the notch filter of the pump with and without going through the device in the linear regime, as well as in the above-OPO regime, is shown in Figure S9.
S13: OSA, ESA and PD measurements of the combs
@@
The detailed pump power and pump wavelength dependence of the topological combs are summarized in Figure S10. The top row shows the measured comb powers as a function of pump wavelength over the 1547.6 nm to 1458.1 nm range of the nonlinear measurements. The edge combs have higher power than the bulk counterparts except for the very high pumping regime. The middle and bottom panels show the corresponding OSA and ESA spectrums. The pump was notched out in all of these measurements.
S14: Quadratic fitting of the comb optical linewidth
@@
We analyzed the optical linewidth of nested comb lines for a selected single ring mode . The results for the nested teeth 1 to 7 are shown in Figure S11 for several pump wavelengths. We clearly observe a quadratic form (with some variation) of the optical linewidth, which goes close to the pump’s linewidth near the pumped mode for most cases. Moreover, the tooth with the minimum linewidth follows the pump wavelength sweeping direction. This behavior is only observed for the edge combs, and the bulk counterparts lack such features.
@@
Moreover, Figure S12 shows the quadratic variation of the comb tooth optical linewidth over both fast and slow timescales for an edge comb. This serves as direct evidence for simultaneous mode-locking over both timescales.
@@
Further, we compare the linewidth of our topological combs with a comparable (with an identical design) single-ring race-track counterpart. Figure S13a shows a generated comb using the same pump laser in our experimental setup. As shown in Figure S13b, the optical linewidth of the comb teeth lacks the quadratic trend of the topological counterpart and also exhibits a much broader linewidth compared to the pump’s linewidth. Moreover, as shown in Figure S13a, nested structure is absent in the comb teeth of the single ring comb. We note that the absence of a quadratic trend across the single-ring comb optical linewidth can be in principle further substantiated by performing the linewidth analysis over a broader range of comb teeth. However, the low SNR of the comb lines away from the pump limits our ability to perform reliable fitting analysis.
S15: Temporal analysis of the pump power dependence of the bulk combs
@@
For a typical bulk comb, the temporal dynamic of the comb is measured as a function of pump power. The results are shown in Figure S14. It can be seen that no appreciable temporal patterns emerge above the OPO threshold, leading to no appearance of a well-defined frequency component in the (Fourier transform) FT analysis.
S16: Temporal analysis of the pump power dependence of the edge combs
@@
For a typical edge comb, the temporal dynamics of the comb are measured as a function of pump power. The results are shown in Figure S15. Unlike the bulk combs, it can be seen that temporal patterns emerge above the OPO threshold, leading to the appearance of well-defined frequency components in the FT analysis.
@@
S17: Temporal analysis of the pump power and pump wavelength dependence of the edge and bulk combs
@@
For several edge and bulk combs, the temporal dynamics of the comb are measured as a function of pump power. The results are shown in Figure S16. The sharp contrast between edge combs and their bulk counterparts can be seen.
S18: Temporal analysis of the combs as a function of pump wavelength.
@@
For a fixed pump power of 185 mW, the temporal dynamics of the combs are measured as a function of pump wavelength. The results are shown in Figure S17. The gradual emergence of temporal patterns as the pump is tuned towards the edge band can be seen.
S19: Transient dynamics of the lattice
Due to the pulsed nature of our laser, one important question is whether the transient behavior, namely the response time of the system until it reaches steady state, is comparable to the 5 ns pulse duration. Figure S18 shows the transient dynamics of our system from both experiment and simulation. It can be seen that from both the experiment and simulation, the duration of transient behavior that is dominated by damped oscillations is about 400 ps, which is at least an order of magnitude lower than the 5 ns pulse duration. The physical reason is that both the intrinsic and extrinsic losses of individual rings are high: the loaded quality factor for a single ring is measured to be only a few thousand.
S20: Filtered-comb experiment
The signatures of mode locking have been demonstrated from both time and frequency domain measurements in the main text. However, several limitations are posed by our laser. Firstly, the finite pulse duration sets a lower limit for frequency-domain resolution, causing our ESA beat notes to be broader than 223 MHz. Secondly, the broad laser background in the frequency domain diminishes the quality of imaging and adds substantial noise in temporal measurements. To address these issues, we split the comb with 1600 nm long pass and short pass filters, and perform the measurements as described in the main text on each part of the comb. Additionally, we use another device with the same design parameters to show the reproducibility of the results demonstrated in the main text.
Figure S20 shows the OSA and ESA measurements for combs with and without filters. The OSA measurements clearly demonstrate the effect of the filters we use, where the pump background is included in the 1600 nm short pass spectrum, but completely filtered out for the 1600 nm long pass spectrum. For all filtering schemes, we see the same ESA beat notes around 4 GHz with near-transform-limit linewidth, corresponding to the slow-time repetition rate.
Next, we perform imaging of the system with the 1580 nm and 1600 nm long pass filter. The pump background is completely filtered out with the 1600 nm filter, but remains for 1580 nm filter. The results are shown in Figure S19. We notice a considerable improvement in the image when we filter out the complete pump background.
We then generate combs with a fixed wavelength in the edge band, and proceed to analyze both filtered parts of the comb in the time domain with the fast (20 GHz) oscilloscope, as shown in Figure S21. By removing the pump background with a 1600 nm long pass filter, and contrasting the result against the 1600 nm short pass filter, where the pump background remains, we have greatly improved the signal to noise ratio of the measurement, as shown in panel a. Remarkably, by fitting the temporal pulses, we show that these pulses can go as narrow as 50 ps, which is a limit set by the oscilloscope (20 GHz). To confirm that these pulses do not come from transient behavior, we fix the wavelength and pump the system below OPO threshold. The measured temporal dynamics in panel b indeed shows that the comb pulses do not come from transient behavior.
Similar to Sections S16 and S17, we have performed sweeps on both pump power (Figure S22) and pump wavelength (Figure S23), showing again the sharp contrast between bulk combs and edge combs, as well as the power dependence of the edge combs in time domain.
To simultaneously show the temporal signatures of mode locking for both parts of the comb (above and below 1600 nm), we have added a supplementary video 1. The video sweeps over 200 independent experiments, showing not only the correlation between two parts of comb in time, but also the repeatability of the results.
S21: Spacing between oscillating edge modes
To better demonstrate the spacing between oscillating edge modes demonstrated in Figure 2d, which is approximately 20 pm, we plot the optical spectrum when we pump in the edge with wavelength 1547.78 nm, as shown in Figure S24. Here, we choose two single-ring modes , and clearly shows that the spacing between super-ring modes is about 20 pm.
We note that, due to the ultra-high resolution of the OSA (0.04 pm), the transmission spectrum has a very high sampling rate even within a narrow 0.2 nm window. Moreover, since the damage threshold of the grating-based OSA is much lower than conventional grating-based counterparts, we attenuated the input optical signal prior to the OSA, which leads to a low signal-to-noise ratio. These two factors combined give the comb teeth spectrum a slightly noisy and spiky appearance.
S22: Limitations of the experiment
In this study, we identify strong signatures of mode locking with various independent measurements. However, direct phase measurements to show mode locking are not presented, and this is due to several experimental challenges and limitations in the current stage. A detailed breakdown of the challenges are as follows:
Coupling 261 rings together. The topological ring lattice contains hundreds of single rings leading to huge fabrication variations. One has to intentionally design a very low quality factor for individual rings (a few thousand) so that the whole device works.
High OPO threshold. A very low quality factor leads to a very high OPO threshold. In our experiment, the peak power of the pulse has to reach about 100W on threshold. On the other hand, using a CW laser with 100W introduces detrimental heat effects that introduce instability and damage chips. Consequently, we use a pulsed laser with a strong peak power, but manageable average power.
Limited pulse duration. In the experiment, the laser duration is 5 ns, which is more than three orders of magnitude longer than single-ring round trip time, at least an order of magnitude longer than both the super-ring round-trip time and the transient time of the system. Even though such a quasi-CW laser is enough to reveal beat notes on frequency measurements, its limited duration sets a lower resolution in frequency domain, as shown in our ESA measurement.
No CW laser. The pulsed nature of the pump stops us from continuously tuning the pump wavelength, which is a common way to find solitons, and is a proposed scheme to find topological solitons in (?). Note that we are not working in the same parameter regime as that described in (?), therefore we are not claiming the existence of solitons based on this study.
Laser jitter. The laser used in this study jitters in time, as shown in Figure 4a, where 2000 independent measurements are plotted on top of each other. The jitter of the laser adds to the broadening of the ESA beat note.
Laser background. The laser used in this study has a spectral background that introduces scattering in imaging and adds noise to both frequency and time-domain measurements. To address this issue, we use spectral filters to eliminate the effect of backgrounds.
To provide stronger evidence of mode locking and its type (Turing rolls, topological solitons), much improvement is needed for the experiment. We note that going forward, an important direction is to improve the Q factor of the system to enable narrow-linewidth CW excitation while reducing the lattice size to make sure that all individual rings are properly coupled, which can make the metrology-type applications, as well as the generation of topological solitons more accessible.
Movie S1. Simultaneous temporal measurement of different parts of the comb. The attached movie S1 shows 200 independent temporal measurements of the comb with the fast (20 GHz) oscilloscope as we pump the edge band and see signatures of mode locking. On each snapshot, we simultaneously measure two parts of topological edge combs using a 1580 nm spectral filter, where the 1580 nm short pass contains parts of the laser background and the 1580 nm long pass contains only the generated comb teeth. It can be seen from the movie that the two parts of the comb are synchronized in time, both giving temporal pulses at the same moment.